Madsen Motors's bonds have 7 years remaining to maturity. Interest is paid annually; they have a $1,000 par value; the coupon interest rate is 9%; and the yield to maturity is 5%. What is the bond's current market price? Round your answer to the nearest cent.

$
Answer #1

M = $1000, n = 7, i = 5%, C = 9% * $1000 = $90

P = 90 * \frac{1-\frac{1}{(1 + 0.05)^7}}{0.05} + \frac{1000}{(1 + 0.05)^7}

P = $520.77 + $710.68

P = $1,231.5